Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2023 logbook, Contact: F.Wolff-Fabris, A.Galler, L.Samoylova, H.Sinn  Not logged in ELOG logo
Message ID: 159     Entry time: 27 Apr 2023, 00:07
Author: Chan Kim 
Group: SPB 
Subject: Shift summary 

SPB/SFX User Proposal 3333 Shift 2 Summary

X-ray delivery

  • 9.3 keV, 1.7 mJ, 202 pulses @ 564 kHz

  • 2 pulse trial @ 2.8 kHz (2.2MHz / 800). Doubled data rate for high viscocity jet. Seemed to work quite well although not expected to be a regular use mode!

Optical laser delivery

  • Only 819 nm ns laser for the jet visualization

Achievements/Observations

  • Data calibration issues solved (thanks to DRC / ITDM experts)

  • Silicon wafer post sample attenuator was dismounted. elog: 45

  • High viscosity jet delivery @ 10 Hz and 2.8 kHz (2 pulses / train @ max spacing)

  • Photosystem-II (@ 41 mm AGIPD Z motor value, adaptive gain) elog: 85,86, various sample preparations

Issues

  • ONDA restarts required

Instrument status

  • Sample chamber and AGIPD turbos OFF, rough pumping ON

  • shutters etc. closed

To do for next shift

  • Catcher and shroud (inc windows) could do with a clean before any alignment (PEG delivery)

  • Proceed with PCS users

ELOG V3.1.4-7c3fd00